# Inertial antiferromagnetic resonance driven by spin-orbit torques

### v1

@@ V Analyses of polarization and handedness
Finally, the field-like SOT [the $\beta$ term in Eq. (3)] has a negligible impact on the resonance modes. In the peak-height and phase-difference expressions, $\beta$ enters only through the factors $1+\alpha\beta$ and $1+\beta^{2}$, which satisfy $\alpha\beta\ll 1$ and $\beta^{2}\ll 1$ for typical torques induced by the spin Hall effect [42]. Numerical simulations confirm the accuracy of the resulting analytical approximation.
VI Conclusion
In this paper, we investigate resonant magnetization precession and nutation in a bilayer system composed of a uniaxial AFM and an adjacent HM layer, driven by two in-plane, mutually perpendicular alternating currents. The amplitude ratio and phase difference of these currents define the handedness and degree of polarization of the driving SOTs.
Using zero-field AFM resonant spectra and approximate analytical expressions, we show that the resonant frequencies are intrinsic, while the resonant peak heights and polarizations are sensitive to both intrinsic parameters (such as the damping factor and inertial relaxation time) and the polarization of the driving force. We further demonstrate that both nutational and precessional resonant modes always appear in pairs for arbitrary amplitude ratios of the currents – unlike in FMs, where only one nutational and one precessional mode appear when the current amplitudes are equal.
Moreover, the resonant modes in AFMs can continuously evolve through elliptical, circular, and linear polarization states by varying the amplitude ratio of the driving currents. In contrast, the resonant modes in FMs remain nearly circularly polarized. Most notably, we identify a critical current ratio at which the resonant mode becomes linearly polarized and the handedness switches across this point. This critical ratio depends on the spin inertia and the type of resonance and has no counterpart in FMs.
@@ VI Conclusion
Our findings highlight the richer landscape of polarization and handedness in AFM resonant dynamics and suggest potential applications for AFM modes as carriers of information in spintronic devices. This work advances the fundamental understanding of inertial resonant modes and provides a basis for chirality control in spin-wave-based information processing.
VII Acknowledgments
@@
Appendix A steady-state linear modes in FMs driven by SOTs
@@
where $\omega_{K}=\gamma_{0}H_{K}$, with $H_{K}$ being effective anisotropy field defined by $H_{K}=2K_{u}/M_{s}-\mu_{0}M_{s}$, with $K_{u}$ being the magnetocrystalline anisotropy constant, $\mu_{0}$ the vacuum susceptibility, and $M_{s}$ the saturation magnetization. The SOTs $\mathbf{\tau}$ read,
$\mathbf{\tau}=-\rho\left\{\mathbf{m}\times\left[\mathbf{m}\times\left(\mathbf{e}_{z}\times\mathbf{j}_{e}\right)\right]+\beta\mathbf{m}\times\left(\mathbf{e}_{z}\times\mathbf{j}_{e}\right)\right\},$ (31)
@@
$\left(\begin{array}[]{c}m_{x}(t)\\ m_{y}(t)\end{array}\right)=\chi\left(\begin{array}[]{c}j_{x}(t)\\ j_{y}(t)\end{array}\right)$ (32)
@@
$m_{x(y)}=\mathcal{A}_{x(y)}\cos\left(\omega t+\phi_{x(y)}\right).$ (35)
@@
$\mathcal{A}_{x(y)}$ $=$ $\rho\sqrt{\frac{P_{1}J^{2}_{x(y)}+P_{2}J^{2}_{y(x)}+P_{3}J_{x}J_{y}}{\widetilde{\Delta}}},$ (36)
@@
where $\tan^{-1}(x,y)$ denotes the two-argument arctangent, which correctly identifies the quadrant of the point $(x,y)$ [47]. The frequency-dependent parameters $Q_{i}$ are defined as $Q_{1}=\Omega[(1+2\alpha\beta-\alpha^{2})\omega^{2}-\Omega^{2}]$, $Q_{2}=\omega[(1+\alpha^{2})(1+\alpha\beta)\omega^{2}-(1-\alpha\beta)\Omega^{2}]$, $Q_{3}=\omega[(1+\alpha^{2})(\alpha-\beta)\omega^{2}+(\alpha+\beta)\Omega^{2}]$, and $Q_{4}=\Omega[(2\alpha-\beta+\alpha^{2}\beta)\omega^{2}+\beta\Omega^{2}]$. From Eqs. (38) and (39), the phase difference between $m_{x}$ and $m_{y}$ is calculated as
@@
Figure 7: (color online). Resonant spectra for various values of $\eta$. Peaks enclosed by circles (ellipses) correspond to precessional (nutational) resonances. To enhance visibility, both axes are shown on a logarithmic scale, and low-frequency regions are omitted. All spectra are calculated from Eq. (36) using the following parameters: $\omega_{K}=0.14$ THz, $\alpha=0.01$, $\beta=0.02$, $\rho\approx 0.13$ Hz///(A///m2); in (a)-(d), $J_{x}=1$ GA///m2, $J_{y}=2$ GA///m2; in (e) and (f), $J_{x}=J_{y}=1$ GA///m2. For comparison with AFMs, the values of $\omega_{K}$ and $\rho$ are chosen to match those in Fig. 2.
Appendix B resonant modes in FMs driven by SOTs
@@
Without spin inertia ($\eta=0$), the resonant frequency is ±$\pm\omega_{K}$. In the presence of spin inertia, the resonant frequencies of nutation ±$\pm\omega_{n}$ and precession ±$\pm\omega_{p}$ are given by
@@
Adopting the same approximation and substituting $\omega$ into Eq. (40), the phase differences corresponding to the ±$\pm\omega_{n}$ and ±$\pm\omega_{p}$ resonant modes are given by
@@
Eqs. (43) and (44) imply that, for the circularly polarized SOTs ($J_{x}=J_{y}$), the nutational mode at $\omega_{n}$ and the precessional mode at $-\omega_{p}$ are nearly suppressed, as shown in Figs. 7(e) and 7(f). Meanwhile, Eqs. (45) and (46) indicate that the phase differences corresponding to the $-\omega_{n}$ and $\omega_{p}$ resonances are both $90^{\circ}$. Consequently, one finds $-\omega_{n}\sin\Delta\phi_{n}^{-}<0$ and $\omega_{p}\sin\Delta\phi_{p}^{+}>0$, indicating that only the lefthanded nutational mode and righthanded precessional mode are excited by circularly polarized SOTs. This is also illustrated in the third column of Fig. 8. These results qualitatively agree with Ref. 29, where a circularly polarized oscillating magnetic field was used to drive the system.
However, for elliptically polarized SOTs ($J_{x}\neq J_{y}$), all resonant modes – both at positive and negative frequencies – can be excited, as shown in Figs. 7(a)-(d). Moreover, all modes are nearly circularly polarized. For nutational resonances, the phase differences satisfy $180^{\circ}<\Delta\phi_{n}^{+}<360^{\circ}$ and $0^{\circ}<\Delta\phi_{n}^{-}<180^{\circ}$, leading to ±$\pm\omega_{n}\sin\Delta\phi_{n}^{\pm}<0$. Therefore, both the positive- and negative-frequency nutational modes (first and third rows of Fig. 8, respectively) are lefthanded. Similarly, for precessional resonances, one has $0^{\circ}<\Delta\phi_{p}^{+}<180^{\circ}$ and $180^{\circ}<\Delta\phi_{p}^{-}<360^{\circ}$, which results in ±$\pm\omega_{p}\sin\Delta\phi_{p}^{\pm}>0$. Consequently, both the positive- and negative-frequency precessional modes (second and fourth rows of Fig. 8) are righthanded. In summary, regardless of whether the SOT is righthanded or lefthanded, nutational resonant modes in FMs are always lefthanded, while precessional resonant modes are always righthanded.
Figure 8: (color online). Polarization and handedness of resonant modes for positive resonant frequencies and various ratios $J_{y}/J_{x}$. The inertial relaxation time is set to $\eta=100$ fs. In each panel, the horizontal and vertical axes represent the $m_{x}$ and $m_{y}$ components, respectively. The axis ranges are scaled by the larger of these two components. The curves depict the trajectories of the magnetization vector, and the arrows indicate the sense of rotation.

### v2

@@ V Analyses of polarization and handedness
Finally, the fieldlike SOT [the $\beta$ term in Eq. (3)] has a negligible impact on the resonance modes. In the peak-height and phase-difference expressions, $\beta$ enters only through the factors $1+\alpha\beta$ and $1+\beta^{2}$, and the contribution of the $\beta$ term is small due to $\alpha\beta\ll 1$ and $\beta^{2}\ll 1$ for typical torques induced by the spin Hall effect [AManchon]. Numerical simulations confirm the accuracy of the analytical approximation.
VI Discussions
First, it is worthwhile to elucidate the significance of the negative frequency. When solving the secular equation for a linear oscillating system, the eigenfrequencies usually come in a pair of opposite numbers. For natural oscillation of magnetization, there is no difference between the positive and negative frequency branches, apart from the opposite handedness of oscillating mode. Thus, the negative eigenfrequency has often been thought to be unnecessary. However, the case is somewhat complicated for resonance of magnetization. As shown in Figs. 2 and 7, the peak heights are different at positive and negative resonant frequencies. Also, Fig. 5 indicates that the polarization and handedness of resonant modes are different for positive and negative resonant frequencies. The reason is that the frequency of forced oscillation belongs to the driving force. Different signs of frequency correspond to different handednesses of driving force. As illustrated in above sections, the properties of resonant mode depend on the handedness of driving force, resulting in the asymmetry between positive and negative resonant frequencies.
Second, AFMs host both lefthanded and righthanded spin waves [SMRezende], enabling storing information on the handedness (charity) of polarization. This feature has been proposed to design transistors [RCheng, TYu] and logic gates [WYu, CJia] based on spin waves. Thus, it is crucial to detect and manipulate the handedness of spin waves, which has been achieved by inverse spin Hall effect in synthetic AFMs [YShiota]. With regard to the spin waves with zero wave vector, it has been shown that the polarization of resonant mode can be identified and manipulated by using different types of microwave magnetic fields [XChen]. Here, by steering the degree of polarization ($J_{y}/J_{x}$) and the handedness of the alternating SOTs, we realize the switching of polarization and the reversal of handedness for the nutational and precessional resonant AFM modes. The critical value of $J_{y}/J_{x}$ can be tuned by the inertial relaxation time $\eta$, as indicated by Eq. (23) and Fig. 6. On the other hand, by varying $J_{y}/J_{x}$ and observing the abrupt change of the polarization, one can measure the critical values $J_{y}/J_{x}=r_{n,p}$, and then estimate $\eta$, which, so far, is not experimentally determined for AFMs.
Third, it should be pointed out that we only explore the linear dynamics of AFMs driven by alternating SOTs. The results are suitable for weak driving forces. With the drives becoming stronger, some high-order harmonic magnetic responses gradually emerge. Thus, beyond the linear response described in Eq. (5), $m_{k}^{x,y}$ should be expanded into a series of harmonics, including the higher-order terms of $j_{x,y}$. Inserting this ansatz and its time derivative into the ILLG equations, and collecting terms at each harmonic ($e^{-ni\omega t},n=1,2,3,\cdots$), one arrives at the scalar equations about higher-order susceptibilities by matching the coefficients for every harmonic orders. In general, it is complicated to get explicit analytic solutions for the susceptibilities and some numeric treatments must be resorted to. Moreover, we use a minimal model of AFMs, in which a perpendicular uniaxial anisotropy is considered. For more complex geometries – for example, involving the biaxial anisotropy, the Dzyaloshinskii-Moriya interaction, triangular spin textures – the linearization technique can also be used. For noncollinear ground states (e.g., canted AFMs [IBoventer]), fluctuations must first be defined in local frames attached to each sublattice, and the equations of motion are then linearized about those frames. Also, by use of the eigenvalues and eigenvectors of linearized system, one can explore the dispersion and polarization of inertial resonant modes. A more thorough understanding of the higher harmonics and further exploration on the complex AFMs are beyond the scope of this work and will be pursued in the future.
Finally, it is interesting to compare the roles of SOTs and spin-transfer torques in driving magnetic resonances. In SOT-based devices, two mutually perpendicular in-plane charge currents can be applied and tuned independently, generating two independent ac driving torques, analogous to two orthogonal in-plane ac magnetic fields. By adjusting the magnitudes and directions (and, if applicable, phases) of these currents, the total driving torque can be engineered to realize different polarizations and handedness. In contrast, in spin-transfer-torque structures the perpendicular current provides essentially a single driving torque with a fixed direction, which does not allow independent control of the polarization and handedness of the excitation.
VII Conclusion
In this paper, we investigate resonant magnetization precession and nutation in a bilayer system composed of a uniaxial AFM and an adjacent HM layer, driven by two in-plane, mutually perpendicular alternating currents. We find that, contrary to the common view that the handedness of a resonant mode is essentially fixed by the material, the polarization state and handedness of both precessional and nutational AFM modes can be controlled by the alternating SOTs generated by in-plane $J_{x}$ and $J_{y}$ currents. As the amplitude ratio $J_{y}/J_{x}$ of orthogonal driving currents is varied, AFM resonant modes evolve continuously from elliptical to circular and, at critical ratios $r_{n,p}$, to linear polarization. We derive the critical ratios $r_{n,p}$ for the nutational ($n$) and precessional ($p$) modes at which a mode becomes linearly polarized and the mode handedness reverse. Moreover, a given mode can be linear on one AFM sublattice while remaining elliptical on the other.
We find that the critical ratios depend explicitly on the inertial relaxation time $\eta$ and differ between the nutational and precessional resonances, enabling inertia-controlled chirality switching. The polarization and handedness reversals at $r_{n,p}$ provide a practical method to determine inertial relaxation time in SOT-driven AFM resonance experiments.
This tunability is not available in the FM case, where the resonant modes remain nearly circular. A detailed comparison with the FM ILLG case shows that in FMs the nutational mode is always lefthanded and the precessional mode always righthanded, and both remain nearly circular regardless of the drive ellipticity. In AFMs, by contrast, both ellipticity and handedness of the modes depend sensitively on $J_{y}/J_{x}$, $\eta$, and the sign of $\omega$. Both nutational and precessional modes appear in pairs at $\pm\omega_{n,p}$ unlike the FM case where certain branches are strongly suppressed for circular driving.
@@ VII Conclusion
Our findings highlight the richer landscape of polarization and handedness in AFM resonant dynamics and suggest potential applications for AFM modes as carriers of information in spintronic devices. This work advances the fundamental understanding of inertial resonant modes and provides a basis for chirality control in spin-wave-based information processing. We expect that the control of the nutational-mode handedness permits a robust manipulation of nutation magnons in AFMs.
VIII Acknowledgments
@@
Appendix A Linearization of AFM system
To linearize the coupled two-lattices AFM system, it is convenient to expand Eq. (1) in the component form
$\frac{dm_{k}^{x}}{dt}$ $=$ $\omega_{E}\left(m_{k}^{y}m_{3-k}^{z}-m_{k}^{z}m_{3-k}^{y}\right)-\omega_{K}m_{k}^{y}m_{k}^{z}+\alpha\left(m_{k}^{y}\frac{dm_{k}^{z}}{dt}-m_{k}^{z}\frac{dm_{k}^{y}}{dt}\right)+\eta\left(m_{k}^{y}\frac{d^{2}m_{k}^{z}}{dt^{2}}-m_{k}^{z}\frac{d^{2}m_{k}^{y}}{dt^{2}}\right)$ (29)
$-\rho\left[j_{x}m_{k}^{x}m_{k}^{y}+j_{y}\left(m_{k}^{y2}+m_{k}^{z2}\right)\right]+\rho\beta j_{x}m_{k}^{z},$
$\frac{dm_{k}^{y}}{dt}$ $=$ $\omega_{E}\left(m_{k}^{z}m_{3-k}^{x}-m_{k}^{x}m_{3-k}^{z}\right)-\omega_{K}m_{k}^{z}m_{k}^{x}+\alpha\left(m_{k}^{z}\frac{dm_{k}^{x}}{dt}-m_{k}^{x}\frac{dm_{k}^{z}}{dt}\right)+\eta\left(m_{k}^{z}\frac{d^{2}m_{k}^{x}}{dt^{2}}-m_{k}^{x}\frac{d^{2}m_{k}^{z}}{dt^{2}}\right)$ (30)
$+\rho\left[j_{y}m_{k}^{x}m_{k}^{y}+j_{x}\left(m_{k}^{z2}+m_{k}^{x2}\right)\right]+\rho\beta j_{y}m_{k}^{z},$
$\frac{dm_{k}^{z}}{dt}$ $=$ $\omega_{E}\left(m_{k}^{x}m_{3-k}^{y}-m_{k}^{y}m_{3-k}^{x}\right)+\alpha\left(m_{k}^{x}\frac{dm_{k}^{y}}{dt}-m_{k}^{y}\frac{dm_{k}^{x}}{dt}\right)+\eta\left(m_{k}^{z}\frac{d^{2}m_{k}^{y}}{dt^{2}}-m_{k}^{y}\frac{d^{2}m_{k}^{x}}{dt^{2}}\right)$ (31)
$-\rho\left(j_{x}m_{k}^{y}m_{k}^{z}-j_{y}m_{k}^{z}m_{k}^{x}\right)-\rho\beta\left(j_{x}m_{k}^{x}+j_{y}m_{k}^{y}\right),$
where $k=1,2$, denoting the two sublattices. In the main text, it has been assumed that the static equilibrium magnetization lies along the $z$ direction. Thus, for small deviations ($m_{k}^{x}$ and $m_{k}^{y}$) from equilibrium, $m_{k}^{z}$ is unchanged to first-order in small quantities ($m_{1}^{z}\approx 1$ and $m_{2}^{z}\approx-1$). For resonance, the excitation sources $j_{x,y}$ are also small quantities. Then, keeping only terms linear in $m_{k}^{x}$, $m_{k}^{y}$, $j_{x}$, and $j_{y}$ in Eqs. (29)-(31), one can obtain the linear differential equations about $m_{1}^{x}$, $m_{1}^{y}$, $m_{2}^{x}$, and $m_{2}^{y}$, expressed in the matrix form as
$\left(\begin{array}[]{cccc}1&\alpha&0&0\\ -\alpha&1&0&0\\ 0&0&1&-\alpha\\ 0&0&\alpha&1\end{array}\right)\frac{d}{dt}\left(\begin{array}[]{c}m_{1}^{x}\\ m_{1}^{y}\\ m_{2}^{x}\\ m_{2}^{y}\end{array}\right)=\left(\begin{array}[]{cccc}0&-\omega_{E}-\omega_{K}&0&-\omega_{E}\\ \omega_{E}+\omega_{K}&0&\omega_{E}&0\\ 0&\omega_{E}&0&\omega_{E}+\omega_{K}\\ -\omega_{E}&0&-\omega_{E}-\omega_{K}&0\end{array}\right)\left(\begin{array}[]{c}m_{1}^{x}\\ m_{1}^{y}\\ m_{2}^{x}\\ m_{2}^{y}\end{array}\right)+\rho\left(\begin{array}[]{cc}\beta&-1\\ 1&\beta\\ -\beta&-1\\ 1&-\beta\end{array}\right)\left(\begin{array}[]{c}j_{x}\\ j_{y}\end{array}\right).$ (32)
Given that the presumed solutions and the applied currents have a form $\sim e^{-i\omega t}$, Eq. (5) can be obtained from Eq. (32) by computing the inverse of matrix.
Appendix B steady-state linear modes in FMs driven by SOTs
@@ Appendix A Linearization of AFM system
where $\omega_{K}=\gamma_{0}H_{K}$, with $H_{K}$ being effective anisotropy field defined by $H_{K}=2K_{u}/M_{s}-\mu_{0}M_{s}$, with $K_{u}$ being the magnetocrystalline anisotropy constant, $\mu_{0}$ the vacuum susceptibility, and $M_{s}$ the saturation magnetization. The SOTs $\mathbf{\boldsymbol{\tau}}$ read,
$\mathbf{\boldsymbol{\tau}}=-\rho\left\{\mathbf{m}\times\left[\mathbf{m}\times\left(\mathbf{e}_{z}\times\mathbf{j}_{e}\right)\right]+\beta\mathbf{m}\times\left(\mathbf{e}_{z}\times\mathbf{j}_{e}\right)\right\},$ (35)
@@
$\left[\begin{array}[]{c}m_{x}(t)\\ m_{y}(t)\end{array}\right]=\chi\left[\begin{array}[]{c}j_{x}(t)\\ j_{y}(t)\end{array}\right]$ (36)
@@
$m_{x(y)}=\mathcal{A}_{x(y)}\cos\left[\omega t+\phi_{x(y)}\right].$ (39)
@@
$\mathcal{A}_{x(y)}$ $=$ $\rho\sqrt{\frac{P_{1}J_{x(y)}^{2}+P_{2}J_{y(x)}^{2}+P_{3}J_{x}J_{y}}{\widetilde{\Delta}}}$, (40)
@@
where $\tan^{-1}(x,y)$ denotes the two-argument arctangent, which correctly identifies the quadrant of the point $(x,y)$ [arctan]. The frequency-dependent parameters $Q_{i}$ are defined as $Q_{1}=\Omega[(1+2\alpha\beta-\alpha^{2})\omega^{2}-\Omega^{2}]$, $Q_{2}=\omega[(1+\alpha^{2})(1+\alpha\beta)\omega^{2}-(1-\alpha\beta)\Omega^{2}]$, $Q_{3}=\omega[(1+\alpha^{2})(\alpha-\beta)\omega^{2}+(\alpha+\beta)\Omega^{2}]$, and $Q_{4}=\Omega[(2\alpha-\beta+\alpha^{2}\beta)\omega^{2}+\beta\Omega^{2}]$. From Eqs. (42) and (43), the phase difference between $m_{x}$ and $m_{y}$ is calculated as
@@
Figure 7: (color online). FM resonant spectra for various values of $\eta$. (a)-(d) Under elliptically polarized SOTs ($J_{x}\neq J_{y}$), both precessional and nutational FM resonances are excited on the positive- and negative-frequency ranges. (e) and (f) For circular polarization ($J_{x}=J_{y}$), the nutational mode is observed at negative frequencies, whereas the precessional mode appears at positive frequencies. Noninertial spectra at $\eta=0$ are shown with dashed curves. The precessional resonance is shifted by finite spin inertia ($\eta>0$) and additional nutational resonances are induced (solid curves). Peaks labeled with red (purple) points correspond to precessional (nutational) resonances. To enhance visibility, both axes are shown on a logarithmic scale, and low-frequency regions are omitted. All spectra are calculated from Eq. (40) using the following parameters: $\omega_{K}=0.14$ THz, $\alpha=0.01$, $\beta=0.02$, $\rho\approx 0.13$ Hz//(A//m2); in (a)-(d), $J_{x}=1$ GA//m2, $J_{y}=2$ GA//m2; in (e) and (f), $J_{x}=J_{y}=1$ GA//m2. For comparison with AFMs, the values of $\omega_{K}$ and $\rho$ are chosen to match those in Fig. 2.
Appendix C resonant modes in FMs driven by SOTs
@@
Without spin inertia ($\eta=0$), the resonant frequency is $\pm\omega_{K}$. In the presence of spin inertia, the resonant frequencies of nutation $\pm\omega_{n}$ and precession $\pm\omega_{p}$ are given by
@@
Adopting the same approximation and substituting $\omega$ into Eq. (44), the phase differences corresponding to the $\pm\omega_{n}$ and $\pm\omega_{p}$ resonant modes are given by
@@
Eqs. (47) and (48) imply that, for the circularly polarized SOTs ($J_{x}=J_{y}$), the nutational mode at $\omega_{n}$ and the precessional mode at $\omega_{p}$ are nearly suppressed, as shown in Figs. 7(e) and 7(f). Meanwhile, Eqs. (49) and (50) indicate that the phase differences corresponding to the $\omega_{n}$ and $\omega_{p}$ resonances are both $90^{\circ}$. Consequently, one finds $\omega_{n}\sin\Delta\phi_{n}^{-}<0$ and $\omega_{p}\sin\Delta\phi_{p}^{+}>0$, indicating that only the lefthanded nutational mode and righthanded precessional mode are excited by circularly polarized SOTs. This is also illustrated in the third column of Fig. 8. These results qualitatively agree with Ref. [RMondal_PRB104_21], where a circularly polarized oscillating magnetic field was used to drive the system.
However, for elliptically polarized SOTs ($J_{x}\neq J_{y}$), all resonant modes – both at positive and negative frequencies – can be excited, as shown in Figs. 7(a)-(d). Moreover, all modes are nearly circularly polarized. For nutational resonances, the phase differences satisfy $180^{\circ}<\Delta\phi_{n}^{+}<360^{\circ}$ and $0^{\circ}<\Delta\phi_{n}^{-}<180^{\circ}$, leading to $\pm\omega_{n}\sin\Delta\phi_{n}^{\pm}<0$. Therefore, both the positive- and negative-frequency nutational modes (first and third rows of Fig. 8, respectively) are lefthanded. Similarly, for precessional resonances, one has $0^{\circ}<\Delta\phi_{p}^{+}<180^{\circ}$ and $180^{\circ}<\Delta\phi_{p}^{-}<360^{\circ}$, which results in $\pm\omega_{p}\sin\Delta\phi_{p}^{\pm}>0$. Consequently, both the positive- and negative-frequency precessional modes (second and fourth rows of Fig. 8) are righthanded. In summary, regardless of whether the SOT is righthanded or lefthanded, nutational resonant modes in FMs are always lefthanded, while precessional resonant modes are always righthanded.
Figure 8: (color online). Polarization and handedness of resonant modes for various ratios $J_{y}/J_{x}$. The inertial relaxation time is set to $\eta=100$ fs. In each panel, the horizontal and vertical axes represent the $m_{x}$ and $m_{y}$ components, respectively. The curves depict the trajectories of the magnetization vector, and the arrows indicate the sense of rotation. In the panels in columns 1-3, $J_{x}=1$GA//m2 and $J_{y}$ is calculated by the ratio $J_{y}/J_{x}$. In the panels in columns 4-5, $J_{y}=1$GA//m2 and $J_{x}$ is calculated by the ratio.